write an equation in slope -intercept form for the line through (-5,6) and (1,0)



answer: y=-x +1
why is the answer y=-x+1 need full break down of why y=-x+1 is an equation in slope -intercept form for the line through (-5,6) and (1,0)

Found 3 solutions by Mathtut, checkley77, solver91311:
Answer by Mathtut(3670)

first find the slope which is change in y over change in x
:
(6-1)/(-5-1)=-1
:
now use point slope formula which is y-k=m(x-h)..where m is the slope and (h,k) is any point on the line. So you can choose either point. I will use (1,0)
:
y-0=-1(x-1)
:
y=-1x+1

Answer by checkley77(12844)

Y=mX+b
(-5,6) and (1,0)
slope(m)=(y2-y1)/(x2-x1)
slope=(0-6)/(1+5)
slope=-6/6
slope=-1
0=-1*1+b
0=-1+b
b=1 the y intercept.
Y=-X+1

Answer by solver91311(24713)



First use the two-point form of the equation of a line:



Where the given points are: and

Substituting the values, given that your points are: and



Do the arithmetic:





Now solve for to put the equation in slope-intercept form, :





Notice that and

John